# Power allocation for cell-free MIMO integrated sensing and communication

### v1

@@ III Sensing Metric
First define the squared effective bandwidth (SEBW) and squared effective pulse time width (SETW) as
@@
[…]
respectively. The cross-term between time and frequency is further defined as
[…]
We further define the average frequency as
[…]
and average time as
[…]
With a relatively large bandwidth and pulse time width, such that $\bar{f}_{n}^{2}\ll\bar{f_{n}^{2}}$ and $(\bar{t}_{n}+\tau_{n,k})^{2}\ll\bar{t_{n}^{2}}$, we have derived the approximate CRLBs for both location and velocity [42], respectively given by
${\bm{C}}_{\rm L}$ $=({\bm{P}}-{\bm{V}}{\bm{Y}}^{-1}{\bm{V}}^{\rm T})^{-1},$ (13)
and
${\bm{C}}_{\rm V}$ $=({\bm{Y}}-{\bm{V}}^{\rm T}{\bm{P}}^{-1}{\bm{V}})^{-1}.$ (14)
The corresponding $2\times 2$ sub-matrix ${\bm{P}}$ is given in (15), ${\bm{V}}$ given in (16),
${\bm{P}}=\sum_{n=1}^{N}\sum_{k=1}^{K}\begin{bmatrix}\beta_{n,k}^{2}a_{n,k}+2\beta_{n,k}\eta_{n,k}b_{n,k}+\eta_{n,k}^{2}d_{n,k}&\beta_{n,k}\zeta_{n,k}a_{n,k}+(\beta_{n,k}\kappa_{n,k}+\zeta_{n,k}\eta_{n,k})b_{n,k}+\eta_{n,k}\kappa_{n,k}d_{n,k}\\ \zeta_{n,k}\beta_{n,k}a_{n,k}+(\zeta_{n,k}\eta_{n,k}+\beta_{n,k}\kappa_{n,k})b_{n,k}+\kappa_{n,k}\eta_{n,k}d_{n,k}&\zeta_{n,k}^{2}a_{n,k}+2\zeta_{n,k}\kappa_{n,k}b_{n,k}+\kappa_{n,k}^{2}d_{n,k}\end{bmatrix},$ (15)
${\bm{V}}=\sum_{n=1}^{N}\sum_{k=1}^{K}\begin{bmatrix}\beta_{n,k}\xi_{n,k}b_{n,k}+\eta_{n,k}\xi_{n,k}d_{n,k}&\beta_{n,k}\varrho_{n,k}b_{n,k}+\eta_{n,k}\varrho_{n,k}d_{n,k}\\ \zeta_{n,k}\xi_{n,k}b_{n,k}+\kappa_{n,k}\xi_{n,k}d_{n,k}&\zeta_{n,k}\varrho_{n,k}b_{n,k}+\kappa_{n,k}\varrho_{n,k}d_{n,k}\end{bmatrix},$ (16)
and
${\bm{Y}}=\sum_{n=1}^{N}\sum_{k=1}^{K}\begin{bmatrix}\xi_{n,k}^{2}d_{n,k}&\xi_{n,k}\varrho_{n,k}d_{n,k}\\ \varrho_{n,k}\xi_{n,k}d_{n,k}&\varrho_{n,k}^{2}d_{n,k}\end{bmatrix}.$ (17)
The involved waveform-related parameters, e.g., $a_{n,k}$, $b_{n,k}$ and $d_{n,k}$, are respectively given by,
$a_{n,k}$ $=\frac{8\pi^{2}|\alpha_{n,k}|^{2}f_{s}P\rho_{n}\bar{f_{n}^{2}}}{\sigma_{z}^{2}},$ (18a)
$b_{n,k}$ $=\frac{4\pi|\alpha_{n,k}|^{2}f_{s}P\rho_{n}}{\sigma_{z}^{2}}\mathfrak{T}\{\sigma_{tf}\},$ (18b)
$d_{n,k}$ $\approx\frac{8\pi^{2}|\alpha_{n,k}|^{2}f_{s}P\rho_{n}\bar{t_{n}^{2}}}{\sigma_{z}^{2}},$ (18c)
where the approximation holds in (18c) when the effective time width meets $\tau_{n,k}\ll T_{\rm eff}$. The relevant geometric spread parameters, including $\beta_{n,k}$, $\zeta_{n,k}$, $\xi_{n,k}$, $\varrho_{n,k}$, $\eta_{n,k}$ and $\kappa_{n,k}$, are omitted here and can be referred to [34, 42].
To visualize the relationship between the CRLBs and the PA coefficients, we now reformulate the CRLB expressions. Given the equivalent location FIM Let $p_{ij}$ denote the $\{i,j\}$th element of matrix ${\bm{P}}$, similar for $v_{ij}$ and $y_{ij}$. The elements $p_{ij}$, $v_{ij}$ and $y_{ij}$ can be expressed as the weighted sum of $\rho_{n}$, i.e.,
$\begin{cases}p_{ij}=\sum_{n=1}^{N}w_{n}^{p_{ij}}\rho_{n}={\bm{\rho}}^{\rm T}{\bm{w}}_{p_{ij}},\\ v_{ij}=\sum_{n=1}^{N}w_{n}^{v_{ij}}\rho_{n}={\bm{\rho}}^{\rm T}{\bm{w}}_{v_{ij}},\\ y_{ij}=\sum_{n=1}^{N}w_{n}^{y_{ij}}\rho_{n}={\bm{\rho}}^{\rm T}{\bm{w}}_{y_{ij}},\end{cases}$ (19)
@@
where $w_{n}^{p_{ij}}$ is the $n$th element of the weight vector ${\bm{w}}_{p_{ij}}$, similar for $w_{n}^{v_{ij}}$ and $w_{n}^{y_{ij}}$. These weights can be readily obtained from (15)-(17) by extracting the PA vector ${\bm{\rho}}$.
By substituting (19) into (13), the CRLB matrix for location estimation can be rewritten as,
${\bm{C}}_{\rm L}=\frac{1}{|{\bm{U}}|}\begin{bmatrix}u_{22}&-u_{12}\\ -u_{21}&u_{11}\end{bmatrix},$ (20)
where $|{\bm{U}}|=u_{11}u_{22}-u_{12}u_{21}$ and
$u_{11}=\frac{1}{|{\bm{Y}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{p_{11}}{\bm{\rho}}^{\rm T}{\bm{W}}_{y}-{\bm{w}}_{v_{11}}{\bm{\rho}}^{\rm T}{\bm{W}}_{12}^{y}-{\bm{w}}_{v_{12}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{11}^{y}){\bm{\rho}}\big{]},$ (21a)
$u_{12}=\frac{1}{|{\bm{Y}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{p_{12}}{\bm{\rho}}^{\rm T}{\bm{W}}_{y}-{\bm{w}}_{v_{21}}{\bm{\rho}}^{\rm T}{\bm{W}}_{12}^{y}-{\bm{w}}_{v_{22}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{11}^{y}){\bm{\rho}}\big{]},$ (21b)
$u_{22}=\frac{1}{|{\bm{Y}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{p_{22}}{\bm{\rho}}^{\rm T}{\bm{W}}_{y}-{\bm{w}}_{v_{21}}{\bm{\rho}}^{\rm T}{\bm{W}}_{22}^{y}-{\bm{w}}_{v_{22}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{21}^{y}){\bm{\rho}}\big{]},$ (21c)
where $|{\bm{Y}}|={\bm{\rho}}^{\rm T}{\bm{W}}_{y}{\bm{\rho}}$. Elements $u_{ij}$ are respectively given by,
$\begin{cases}{\bm{W}}_{y}={\bm{w}}_{y_{11}}{\bm{w}}_{y_{22}}^{\rm T}-{\bm{w}}_{y_{12}}{\bm{w}}_{y_{21}}^{\rm T},\\ {\bm{W}}^{y}_{12}={\bm{w}}_{v_{11}}{\bm{w}}_{y_{22}}^{\rm T}-{\bm{w}}_{v_{12}}{\bm{w}}_{y_{21}}^{\rm T},\\ {\bm{W}}^{y}_{22}={\bm{w}}_{v_{21}}{\bm{w}}_{y_{22}}^{\rm T}-{\bm{w}}_{v_{22}}{\bm{w}}_{y_{21}}^{\rm T},\\ \breve{\bm{W}}^{y}_{11}={\bm{w}}_{v_{12}}{\bm{w}}_{y_{11}}^{\rm T}-{\bm{w}}_{v_{11}}{\bm{w}}_{y_{12}}^{\rm T},\\ \breve{\bm{W}}^{y}_{21}={\bm{w}}_{v_{22}}{\bm{w}}_{y_{11}}^{\rm T}-{\bm{w}}_{v_{21}}{\bm{w}}_{y_{12}}^{\rm T}.\end{cases}$ (22)
Similarly, the velocity CRLB matrix can be rewritten as,
${\bm{C}}_{\rm V}=\frac{1}{|{\bm{H}}|}\begin{bmatrix}h_{22}&-h_{12}\\ -h_{21}&h_{11}\end{bmatrix},$ (23)
where $|{\bm{H}}|=h_{11}h_{22}-h_{12}h_{21}$. Elements $h_{ij}$ are respectively given by,
$h_{11}=\frac{1}{|{\bm{P}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{y_{11}}{\bm{\rho}}^{\rm T}{\bm{W}}_{p}-{\bm{w}}_{v_{11}}{\bm{\rho}}^{\rm T}{\bm{W}}_{12}^{p}-{\bm{w}}_{v_{21}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{11}^{p}){\bm{\rho}}\big{]},$ (24a)
$h_{12}=\frac{1}{|{\bm{P}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{y_{12}}{\bm{\rho}}^{\rm T}{\bm{W}}_{p}-{\bm{w}}_{v_{12}}{\bm{\rho}}^{\rm T}{\bm{W}}_{12}^{p}-{\bm{w}}_{v_{22}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{11}^{p}){\bm{\rho}}\big{]},$ (24b)
$h_{22}=\frac{1}{|{\bm{P}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{y_{22}}{\bm{\rho}}^{\rm T}{\bm{W}}_{p}-{\bm{w}}_{v_{12}}{\bm{\rho}}^{\rm T}{\bm{W}}_{22}^{p}-{\bm{w}}_{v_{22}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{21}^{p}){\bm{\rho}}\big{]},$ (24c)
@@
$\begin{cases}{\bm{W}}_{p}={\bm{w}}_{p_{11}}{\bm{w}}_{p_{22}}^{\rm T}-{\bm{w}}_{p_{12}}{\bm{w}}_{p_{21}}^{\rm T},\\ {\bm{W}}^{p}_{12}={\bm{w}}_{v_{11}}{\bm{w}}_{p_{22}}^{\rm T}-{\bm{w}}_{v_{21}}{\bm{w}}_{p_{21}}^{\rm T},\\ {\bm{W}}^{p}_{22}={\bm{w}}_{v_{12}}{\bm{w}}_{p_{22}}^{\rm T}-{\bm{w}}_{v_{22}}{\bm{w}}_{p_{21}}^{\rm T},\\ \breve{\bm{W}}^{p}_{11}={\bm{w}}_{v_{21}}{\bm{w}}_{p_{11}}^{\rm T}-{\bm{w}}_{v_{11}}{\bm{w}}_{p_{12}}^{\rm T},\\ \breve{\bm{W}}^{p}_{21}={\bm{w}}_{v_{22}}{\bm{w}}_{p_{11}}^{\rm T}-{\bm{w}}_{v_{12}}{\bm{w}}_{p_{12}}^{\rm T}.\end{cases}$ (25)
IV Problem Formulation
In general, a higher power budget leads to improved sensing and communication performance. This section investigates a novel PA scheme aimed at maximizing the overall performance of CF MIMO-ISAC systems. Depending on specific application requirements, the optimization objective can be either sensing or communication performance, while the quality of service (QoS) requirement of the other function is incorporated as a constraint in the optimization model.
Let ${\bm{\rho}}_{\min}=[\rho_{\min,1},\rho_{\min,2},\cdots,\rho_{\min,N}]^{\rm T}$ denote the minimum PA coefficient to satisfy the so-called larger SNR for tight CRLBs, and ${\bm{\rho}}_{\max}=[\rho_{\max,1},\rho_{\max,2},\cdots,\rho_{\max,N}]^{\rm T}$ denote the maximum PA coefficient due to the power budgets of individual transmitters [40, 13]. In this paper, we consider maximizing the communication data rate or SINR under the constraint of the sensing performance, i.e.,
$\arg\min\limits_{{\bm{\rho}}}$ $\ -{\bm{\rho}}^{\rm T}{\bm{g}},$ (26a)
${\rm s.t.}$ ${\bm{\rho}}^{\rm T}{\bm{1}}=1,$ (26b)

### v2

@@ III The Proposed Sensing Metric and Problem Formulation
To begin with, we introduce the following definitions. Define the squared effective bandwidth (SEBW) and squared effective pulse time width (SETW) as
@@
[…]
respectively.
III-A Sensing Metric
Unlike the sensing performance metrics employed in [24, 25, 34, 35, 36], which only consider target detection or location estimation, we consider the joint location and velocity CRLB as the sensing metric, developed based on the discrete-time reception of the reflected signal under the deterministic RCS assumption. In particular, we have derived the approximate CRLBs for both location and velocity in [30] for simplified sensing metrics. They are respectively given by
$\small{\bm{C}}_{\rm L}$ $=({\bm{P}}-{\bm{V}}{\bm{Y}}^{-1}{\bm{V}}^{\rm T})^{-1},$ (11)
@@ III-A Sensing Metric
$\small{\bm{C}}_{\rm V}$ $=({\bm{Y}}-{\bm{V}}^{\rm T}{\bm{P}}^{-1}{\bm{V}})^{-1}.$ (12)
The detailed expression of ${\bm{P}}$ and ${\bm{V}}$ are given at the top of the page, respectively, and
${\bm{P}}=\sum_{n=1}^{N}\sum_{k=1}^{K}\begin{bmatrix}\beta_{n,k}^{2}a_{n,k}+2\beta_{n,k}\eta_{n,k}b_{n,k}+\eta_{n,k}^{2}d_{n,k}&\beta_{n,k}\zeta_{n,k}a_{n,k}\!+\!(\beta_{n,k}\kappa_{n,k}\!+\!\zeta_{n,k}\eta_{n,k})b_{n,k}+\eta_{n,k}\kappa_{n,k}d_{n,k}\\ \zeta_{n,k}\beta_{n,k}a_{n,k}\!+\!(\zeta_{n,k}\eta_{n,k}+\beta_{n,k}\kappa_{n,k})b_{n,k}+\kappa_{n,k}\eta_{n,k}d_{n,k}&\zeta_{n,k}^{2}a_{n,k}+2\zeta_{n,k}\kappa_{n,k}b_{n,k}\!+\!\kappa_{n,k}^{2}d_{n,k}\end{bmatrix},$
$\small{\bm{V}}=\sum_{n=1}^{N}\sum_{k=1}^{K}\begin{bmatrix}\beta_{n,k}\xi_{n,k}b_{n,k}+\eta_{n,k}\xi_{n,k}d_{n,k}&\beta_{n,k}\varrho_{n,k}b_{n,k}+\eta_{n,k}\varrho_{n,k}d_{n,k}\\ \zeta_{n,k}\xi_{n,k}b_{n,k}+\kappa_{n,k}\xi_{n,k}d_{n,k}&\zeta_{n,k}\varrho_{n,k}b_{n,k}+\kappa_{n,k}\varrho_{n,k}d_{n,k}\end{bmatrix},$
$\small{\bm{Y}}=\sum_{n=1}^{N}\sum_{k=1}^{K}\begin{bmatrix}\xi_{n,k}^{2}d_{n,k}&\xi_{n,k}\varrho_{n,k}d_{n,k}\\ \varrho_{n,k}\xi_{n,k}d_{n,k}&\varrho_{n,k}^{2}d_{n,k}\end{bmatrix}.$ (13)
The relevant waveform-related parameters $a_{n,k}$, $b_{n,k}$ and $d_{n,k}$, and geometric-spreading-related factors $\beta_{n,k}$, $\zeta_{n,k}$, $\xi_{n,k}$, $\varrho_{n,k}$, $\eta_{n,k}$, and $\kappa_{n,k}$, are detailed in Appendix A.
As observed from (61), the PA coefficients play a pivotal role in the formulation of the CRLBs. Meanwhile, these coefficients directly affect communication SINR, as described in (7). Unfortunately, the PA optimization relevant to the CRLB is intractable due to the implicit and complex forms of the expressions in (11) and (12). To facilitate analysis and optimization, we reformulate the CRLB expressions into a more explicit and tractable form.
We denote the $(i,j)$th elements of matrices ${\bm{P}}$, ${\bm{V}}$ and ${\bm{Y}}$ as $p_{ij}$, $v_{ij}$ and $y_{ij}$, respectively. Then, one can simplify them as
$\small\begin{cases}p_{ij}=\sum_{n=1}^{N}w_{n}^{p_{ij}}\rho_{n}={\bm{\rho}}^{\rm T}{\bm{w}}_{p_{ij}},\\ v_{ij}=\sum_{n=1}^{N}w_{n}^{v_{ij}}\rho_{n}={\bm{\rho}}^{\rm T}{\bm{w}}_{v_{ij}},\\ y_{ij}=\sum_{n=1}^{N}w_{n}^{y_{ij}}\rho_{n}={\bm{\rho}}^{\rm T}{\bm{w}}_{y_{ij}},\end{cases}$ (14)
@@
where $w_{n}^{p_{ij}}$, $w_{n}^{v_{ij}}$ and $w_{n}^{y_{ij}}$ are the $n$th element of the weight vector ${\bm{w}}_{p_{ij}}$, ${\bm{w}}_{v_{ij}}$ and ${\bm{w}}_{y_{ij}}$, respectively. These weights can be readily obtained from the definitions of ${\bm{P}}$, ${\bm{V}}$ and ${\bm{Y}}$, by extracting their PA coefficient $\rho_{n}$. For example, from (13), we have $w_{n}^{p_{11}}$ as
$\small w_{n}^{p_{11}}=\sum\nolimits_{k=1}^{K}(\beta_{n,k}^{2}\bar{a}_{n,k}+2\beta_{n,k}\eta_{n,k}\bar{b}_{n,k}+\eta_{n,k}^{2}\bar{d}_{n,k}),$ (15)
where the details of $\bar{a}_{n,k}$, $\bar{b}_{n,k}$ and $\bar{d}_{n,k}$ are referred to (61).
By substituting (14) into (11), the CRLB matrix for location estimation can be rewritten as
$\small{\bm{C}}_{\rm L}=\frac{1}{|{\bm{U}}|}\begin{bmatrix}u_{22}&-u_{12}\\ -u_{21}&u_{11}\end{bmatrix},$ (16)
where $|{\bm{U}}|=u_{11}u_{22}-u_{12}u_{21}$. The relevant elements $u_{ij}$ are given respectively by
$\small u_{11}=\frac{1}{|{\bm{Y}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{p_{11}}{\bm{\rho}}^{\rm T}{\bm{W}}_{y}\!-\!{\bm{w}}_{v_{11}}{\bm{\rho}}^{\rm T}{\bm{W}}_{12}^{y}\!-\!{\bm{w}}_{v_{12}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{11}^{y}){\bm{\rho}}\big{]},$ (17)
$u_{12}=\frac{1}{|{\bm{Y}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{p_{12}}{\bm{\rho}}^{\rm T}{\bm{W}}_{y}\!-\!{\bm{w}}_{v_{21}}{\bm{\rho}}^{\rm T}{\bm{W}}_{12}^{y}\!-\!{\bm{w}}_{v_{22}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{11}^{y}){\bm{\rho}}\big{]},$
$u_{22}=\frac{1}{|{\bm{Y}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{p_{22}}{\bm{\rho}}^{\rm T}{\bm{W}}_{y}\!-\!{\bm{w}}_{v_{21}}{\bm{\rho}}^{\rm T}{\bm{W}}_{22}^{y}\!-\!{\bm{w}}_{v_{22}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{21}^{y}){\bm{\rho}}\big{]},$
where $|{\bm{Y}}|={\bm{\rho}}^{\rm T}{\bm{W}}_{y}{\bm{\rho}}$, and
$\small\begin{cases}{\bm{W}}_{y}={\bm{w}}_{y_{11}}{\bm{w}}_{y_{22}}^{\rm T}-{\bm{w}}_{y_{12}}{\bm{w}}_{y_{21}}^{\rm T},\\ {\bm{W}}^{y}_{12}={\bm{w}}_{v_{11}}{\bm{w}}_{y_{22}}^{\rm T}-{\bm{w}}_{v_{12}}{\bm{w}}_{y_{21}}^{\rm T},\\ {\bm{W}}^{y}_{22}={\bm{w}}_{v_{21}}{\bm{w}}_{y_{22}}^{\rm T}-{\bm{w}}_{v_{22}}{\bm{w}}_{y_{21}}^{\rm T},\\ \breve{\bm{W}}^{y}_{11}={\bm{w}}_{v_{12}}{\bm{w}}_{y_{11}}^{\rm T}-{\bm{w}}_{v_{11}}{\bm{w}}_{y_{12}}^{\rm T},\\ \breve{\bm{W}}^{y}_{21}={\bm{w}}_{v_{22}}{\bm{w}}_{y_{11}}^{\rm T}-{\bm{w}}_{v_{21}}{\bm{w}}_{y_{12}}^{\rm T}.\end{cases}$ (18)
Similarly, the velocity CRLB matrix can be rewritten as
$\small{\bm{C}}_{\rm V}=\frac{1}{|{\bm{H}}|}\begin{bmatrix}h_{22}&-h_{12}\\ -h_{21}&h_{11}\end{bmatrix},$ (19)
where $|{\bm{H}}|=h_{11}h_{22}-h_{12}h_{21}$. The relevant elements $h_{ij}$ are given respectively by
$\small h_{11}=\frac{1}{|{\bm{P}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{y_{11}}{\bm{\rho}}^{\rm T}{\bm{W}}_{p}\!-\!{\bm{w}}_{v_{11}}{\bm{\rho}}^{\rm T}{\bm{W}}_{12}^{p}\!-\!{\bm{w}}_{v_{21}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{11}^{p}){\bm{\rho}}\big{]},$ (20)
$h_{12}=\frac{1}{|{\bm{P}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{y_{12}}{\bm{\rho}}^{\rm T}{\bm{W}}_{p}\!-\!{\bm{w}}_{v_{12}}{\bm{\rho}}^{\rm T}{\bm{W}}_{12}^{p}\!-\!{\bm{w}}_{v_{22}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{11}^{p}){\bm{\rho}}\big{]},$
$h_{22}=\frac{1}{|{\bm{P}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{y_{22}}{\bm{\rho}}^{\rm T}{\bm{W}}_{p}\!-\!{\bm{w}}_{v_{12}}{\bm{\rho}}^{\rm T}{\bm{W}}_{22}^{p}\!-\!{\bm{w}}_{v_{22}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{21}^{p}){\bm{\rho}}\big{]},$
@@
$\small\begin{cases}{\bm{W}}_{p}={\bm{w}}_{p_{11}}{\bm{w}}_{p_{22}}^{\rm T}-{\bm{w}}_{p_{12}}{\bm{w}}_{p_{21}}^{\rm T},\\ {\bm{W}}^{p}_{12}={\bm{w}}_{v_{11}}{\bm{w}}_{p_{22}}^{\rm T}-{\bm{w}}_{v_{21}}{\bm{w}}_{p_{21}}^{\rm T},\\ {\bm{W}}^{p}_{22}={\bm{w}}_{v_{12}}{\bm{w}}_{p_{22}}^{\rm T}-{\bm{w}}_{v_{22}}{\bm{w}}_{p_{21}}^{\rm T},\\ \breve{\bm{W}}^{p}_{11}={\bm{w}}_{v_{21}}{\bm{w}}_{p_{11}}^{\rm T}-{\bm{w}}_{v_{11}}{\bm{w}}_{p_{12}}^{\rm T},\\ \breve{\bm{W}}^{p}_{21}={\bm{w}}_{v_{22}}{\bm{w}}_{p_{11}}^{\rm T}-{\bm{w}}_{v_{12}}{\bm{w}}_{p_{12}}^{\rm T}.\end{cases}$ (21)
Note that the newly derived CRLB expressions in (16) and (19) represent explicit functions of the PA vector, thereby serving as effective and tractable sensing performance metrics.
III-B Problem Formulation
This section formulates the PA problem by maximizing the overall performance of CF MIMO-ISAC systems. Depending on specific application requirements, the optimization objective can be either sensing or communication performance, while the requirement of the other functionality is incorporated as a constraint in the optimization model.
Let ${\bm{\rho}}_{\min}=[\rho_{\min,1},\rho_{\min,2},\cdots,\rho_{\min,N}]^{\rm T}$ denote the minimum PA coefficient to achieve the sufficiently large SNR for tight CRLBs [34, 36], and ${\bm{\rho}}_{\max}=[\rho_{\max,1},\rho_{\max,2},\cdots,\rho_{\max,N}]^{\rm T}$ denote the maximum PA coefficient due to the power budgets of individual transmitters [38, 36]. In this paper, we consider the problem of maximizing the communication SINR, subject to constraints on the sensing performance, i.e.,
$\small\arg\min\limits_{{\bm{\rho}}}$ $\ -{\bm{\rho}}^{\rm T}{\bm{g}},$ (22a)
${\rm s.t.}$ ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}=1,$ (22b)